Solve T(n) = 9T(n/3) + n2 using master's theorem
Answer #1

Answer:

T(n) = 9*T(n/3) +n? This is in the form of T(n) = a1 + f(n) so, we can use masters theorem. compare nlogta with f(n) a = 9,



Please let me know if you have any doubts Please upvote this answer. Thanks!!